<>

CBH Gregorian Dome Receiver Log

as of Saturday, 16-Dec-2017 22:07:06 AST

Go to Most Recent Entry

***Note:*** if you were directed here from the Receiver Tsys page because the system temperature was high, please note that when the receiver is covered with a manual radar cover, the tsys will read high, although the receiver is OK. Check with the operator if in doubt.


Name Email Date Comment
lswlwray@naic.eduMonday, February 2, 2004 17:13:01Receiver up at test facility, first Trx measurements show 11-12K flat across band.
ansanolla@naic.eduWednesday, March 3, 2004 16:03:20System installed upstairs on 2/20/2004. It's Tsys is chann A 24 ēk and chann B is 28 ēk.
ansanolla@naic.eduMonday, March 8, 2004 14:19:48The remote cryotemp reading was hooked up to pos # 6 (former LBN one). The diodes order have to be fix because the 70 diode is hooked to the 16 deg stage and viceversa the 16 deg diode is at 70 deg stage.
ansanolla@naic.eduTuesday, March 9, 2004 13:02:25Dewar shimms added to Dewar , no change to focal point. They are 16" high.
ansanolla@naic.eduWednesday, March 17, 2004 13:35:13Diodes D102 & D502 on bias board been replaced by LED's.
lswlwray@naic.eduThursday, June 10, 2004 16:20:12Moving helium lines around in the cable wrap to reduce twistyness... receiver was off for ~5 minutes around 10am... recovered well.
lswlwray@naic.eduTuesday, June 15, 2004 08:25:24June 14: Dual beam radar causes the temp monitor data to go crazy, much more than the other receivers. Something is much more sensitive in this receiver.
ansanolla@naic.eduFriday, July 2, 2004 11:34:52System down due to power shut down (main power line shorted at the rotary joint).
ansanolla@naic.eduSaturday, July 3, 2004 16:33:18System been cooled.
ansanolla@naic.eduTuesday, July 6, 2004 10:47:19System back to normal after the power emegency on Friday. Vac valve was closed this morning.
ansanolla@naic.eduThursday, July 29, 2004 16:13:16Protective shutter on it.It increases the T Sys.
ansanolla@naic.eduFriday, August 13, 2004 14:47:24CBH shutter on .
ansanolla@naic.eduMonday, August 23, 2004 12:55:26Shutter removed.
ansanolla@naic.eduTuesday, August 31, 2004 20:03:44Automatic shutter installed today at AST 11:30.
ansanolla@naic.eduWednesday, September 22, 2004 13:06:26The shutter was formaly installed today , it first try was on August but had problems on it.
ansanolla@naic.eduThursday, September 30, 2004 14:35:28CBH shutter motor shorted out. It was causing that other shutters wont close. So motor removed in order to replace it, shutter closed by hand. Receiver on schedule for 10/8/2004.
ansanolla@naic.eduMonday, October 11, 2004 15:59:11System was down due to power failure on Saturday , was cooled down by BG and Ganesh.
ansanolla@naic.eduWednesday, November 24, 2004 13:16:44Bias & post amp chassis are off , pin pointing an rfi of 60 Mhz produced localy.
ansanolla@naic.eduMonday, November 29, 2004 13:36:22Bias & post amp boxes on again , when units are off they load cal select ckt. preventing them from changing.
ansanolla@naic.eduTuesday, November 30, 2004 14:31:45The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
ansanolla@naic.eduTuesday, December 21, 2004 08:37:23DB 15 & 25 control cables removed from receiver for rfi test by JRV & CR.
ansanolla@naic.eduTuesday, January 4, 2005 15:51:00System down until further notice.
ansanolla@naic.eduThursday, January 13, 2005 12:55:10System been cooled ,cryocomp ED 010 tie up again with system.
ansanolla@naic.eduFriday, January 14, 2005 11:41:45System cooled. Its cryocomp found with a supply preassure of 240 , charged to 270. Keep an eye on it.
ansanolla@naic.eduFriday, January 14, 2005 11:51:07Forgot to mention that the crosshead was replaced due that the one in broke the motor shaft when it was stuck.It was done in 1/13/2005.
ansanolla@naic.eduWednesday, January 19, 2005 16:04:35Last entry not valid. It belongs to the XB reveiver.
ansanolla@naic.eduThursday, January 20, 2005 12:54:34Noise injection board modified ( to avoid oscillations ,in the SBL it was causing a 60 Mhz interference).Also TTL cal driver board modified (10K resistor at input for a low in condition ).
crvcrios@naic.eduMonday, March 21, 2005 15:49:37Mux board bypass removed and the mux board was reconected.
ansanolla@naic.eduTuesday, March 22, 2005 14:23:00The mux board removed again due to suspicius rfi at the LBW receiver. Its mux board is a no modified one.
ansanolla@naic.eduMonday, May 23, 2005 15:38:22Channel B coax cable from Dewar to cabinet was smash.Replaced.
W. Genterwgenter@naic.eduMonday, August 1, 2005 09:40:04Receiver warmed up on July 29, 2005 due to power dip and compressor failure to restart. Recooled the receiver on the same day.
W. Genterwgenter@naic.eduMonday, August 1, 2005 10:04:14Previous entry for SB High receiver.
ansanolla@naic.eduThursday, August 4, 2005 16:17:44Channel A rf cable(5 ft. long) between Dewar 7 post amp bad, replaced with spare one.Also channel B rf cable (12 ft. long) between post amp & RF cabinet was bad.
ansanolla@naic.eduFriday, August 26, 2005 13:07:48Due to power failure on 8/25/2005 system warmed up.
ansanolla@naic.eduTuesday, August 30, 2005 15:44:28System been cooled. 3 vac pumps on it.
ansanolla@naic.eduWednesday, August 31, 2005 15:44:19System back to normal.
W. Genterwgenter@naic.eduFriday, October 14, 2005 09:07:25Last evening the shutter would not fully open or close. The motor (brushes) caused broadband RFI in all the receivers. When the shutters were commanded to close, the motor timed out and the RFI vanished. On the command to open the shutter, the motor ran continusly. The shutter motor power was removed and the shutter closed and secured.
W. Genterwgenter@naic.eduWednesday, July 26, 2006 16:26:39Two PAMTECH Isolators arrived today. Still waiting on the Mac Tech directional couplers.
ansanolla@naic.eduFriday, August 11, 2006 14:21:35The cryo comp ED 014 has a leak on it.CR & VN working on it.
W. Genterwgenter@naic.eduWednesday, August 16, 2006 07:53:55Mac Tech directional couplers have been received.
ansanolla@naic.eduTuesday, December 5, 2006 13:23:56System cryo down in order to use its cryo comp & lines with the XB/SBN receivers.
crvcrios@naic.eduTuesday, December 12, 2006 13:21:12System satrting to cool down with it's original compresor and lines.
ansanolla@naic.eduThursday, December 14, 2006 16:19:09System back to normal.Own cryo comp & lines.
ansanolla@naic.eduWednesday, January 17, 2007 13:18:50Cryo system shut off in order to use its cryo comp (ED 010) for feeding the LBW cryo system ,LBW cryo comp (ED 007) cooling fan failed.
ansanolla@naic.eduFriday, February 2, 2007 11:28:36Cryo comp ED 010 back to it , it was with the LBW cryo system due to an emergency. Cool down started.
ansanolla@naic.eduMonday, February 5, 2007 11:25:56Dewar vac valve closed. System back to normal.
ansanolla@naic.eduTuesday, April 24, 2007 10:45:20System down due to painting project.
ansanolla@naic.eduWednesday, April 25, 2007 13:00:59The system cryo comp ( ED010) was taken down for service.
ansanolla@naic.eduMonday, May 7, 2007 10:23:42New power supplies added to the bias box (15 vdc for bias was replaced by an Acopian 15DB-30 , it was Lamda one ). Also an Acopian 5EB50 was added to supply the 5 vdc for the digital ckts , they were supplied from the Post amp chassis.The power sw for the bias box was replaced for a locking type one to avoid accidental turn off. The Post amp box power sw was also replaced with a locking type to avoid accidental turn off.
CRVcrios@naic.eduWednesday, January 2, 2008 14:25:19Vacuum pump connected at dewar and is pumping.
crvThursday, July 17, 2008 16:46:24CBH cross head was removed to be used on LBW.
crvFriday, July 18, 2008 14:34:01Cryo pump 024 is installed on the dewar and ready to cool on monday july de 21st 2008
crvMonday, July 21, 2008 11:02:08system started cooldown system sounds good using comp.009.
CRVMonday, October 13, 2008 14:21:24On October 5th 2008 Cryo lines 7 and 8 where cleaned and flushed with petroluem ether to remove existing oil in the lines.Lines where purged with Helium to remove residue of ether in the lines. Lines where left doing vacuum.
CRVMonday, October 13, 2008 14:22:22Dewar Vacuum Started on October the 6th 2008
CRV Monday, October 13, 2008 14:27:03On October the 7th 2008 Comp ed15 lowered from the dome to be repaired and used to cooldown CBH.
CRVMonday, October 13, 2008 14:28:22on october the 8th Comp.015 was repaired and ready for testing on the lab.
CRV Monday, October 13, 2008 14:29:59on October 9th Comp.15 testing started on the lab test dewar will be tested until October 10th 2008
CRV Monday, October 13, 2008 14:32:24On October 10th comp.15 removed from the test dewar and raised to the the dome connected to Sbw because we are going to use comp 9 on on CBH.
CRVMonday, October 13, 2008 14:34:45Cross head ED 030 removed from the dewar and a brand new crosshead installed in the dewar. Cryo lines 7 and 8 where charged and vacuum pump removed. Comp 9 connected to the system and reciever cooling .
CRVMonday, October 13, 2008 14:38:10during the weekend the dewar started to slowly increase its temperature . Guarionex and I when to check the reciever and found tha the comp 9 pressure was in 120 PSI. We checkeed for leaks and found that an o-ring on the supply connector of the compressor was damaged (broken) we replaced the o-ring and the leak stoped.
crvWednesday, November 5, 2008 13:18:19System was warmed up on Tuesday oct. 28th due to cryo pump contamination. system was kept warmed until Thursday Oct. 30th due to 430 Mhz transmitter run.
crvWednesday, November 5, 2008 13:22:35On Thursday Oct. 30th the cryo pump was removed from the dewar to verify the typt of contaminant. The contaminant found was water on the first stage displacer. The dewar and cryopump were cleaned. The system stared the cooling down.
Stephen Jensensjensen@naic.eduThursday, December 11, 2008 13:12:36CBH has been warmed up due to moving its compressor to other receivers.
Stephen Jensensjensen@naic.eduThursday, December 11, 2008 16:26:3111 DEC 4:00 PM: 327 cooling down nicely. X-band starting to cool.
GRgrivera@naic.eduThursday, February 5, 2009 11:16:09On Wednesday, Nov. 12, 2008, CBH crosshead was removed and cleaned. It had a lot of moisture (even though this was done at ~295K) and some charcoal residue on the small displacer. The cylinder was honed and cleaned. After bolting the crosshead to the Dewar, it was purged and vacuumed up to about 40 mTorr. CBH was left cooling.
crvThursday, May 28, 2009 17:04:18 We cleaned the helium lines 7 and 8 of the CBH system with ether and flushed them dry with helium and left the line doing vacuum. The vacuum was at 40u when we came down.
crvFriday, June 12, 2009 13:23:55system lines cleaned with ether and vacuumed on 5/28/2009
crvFriday, June 12, 2009 13:24:50system started cooldown on 5/29/2009
crvFriday, June 12, 2009 13:25:34system cooled well and ready to use on June 1st 2009
crvFriday, June 12, 2009 13:27:26System was warmed down on june 10th 2009 to cooldown SBH.
Stephen Jensensjensen@naic.eduThursday, July 2, 2009 15:29:53CBH is warmed up and will only be cooled for a "campaign". The next scheduled cooling is Aug 8,9 & Sep 19,20.
Stephen Jensensjensen@naic.eduWednesday, July 29, 2009 11:27:14CBH appears to be broken. No signals were seen coming into the post amp. Will probably make preparations to bring CBH down.
Stephen Jensensjensen@naic.eduFriday, July 31, 2009 15:01:32The bias box for CBH was not functioning (channel "B"). The bias box from X-Band was used in place of the CBH bias box.
crvThursday, August 6, 2009 15:13:54reciever was vacuumed and is cooled down today 8/6/2009
crvcrios@naic.eduWednesday, February 24, 2010 16:44:25System was warmed up on Feb. 11th, 2010 due to cryo pump exsecive racheting (noise). The System has started cooldown today Feb. 24th, 2010.
crvcrios@naic.eduMonday, March 22, 2010 14:20:43The system was rachweting during the weekend and the operator turned it off . the system started vacc. today.
crvcrios@naic.eduWednesday, April 7, 2010 15:13:12The system started cooldown today. The system wasn't contaminated.
crvcrios@naic.eduWednesday, June 2, 2010 10:45:32The system was turned off on May 27th,2010. The system will be lowered on the week to make space for the BYU proyect.
crvThursday, September 30, 2010 16:36:20The dewar has been opened to by inspected and do maintnance on it before been reinstalled.
crioscrios@naic.eduThursday, April 14, 2011 10:00:34The dewar was reasembled and the system was tested an is operational. The dewar is ready to be installed.
criosWednesday, May 25, 2011 15:29:21The receiver was installed on the dome and started cooldown on 5/17/2011.
Dana W. & Denis Urbaindwhitlow@naic.eduTuesday, July 19, 2011 17:22:55The all-receiver impulsive RFI problem first noted on Sunday July 17, and first really complained about on Monday night July 18, proved due to motor commutator noise. A gear in the drive train of the CBH shutter mechanism had come loose so that the shutter did not move to its limit switch, leaving the motor running continuously. Lacking access to proper tools, the interim solution was to pull the plug to that motor. NOTE: We must not do any radar from the dome until this is properly fixed!
Dana Whitlowdwhitlow@naic.eduWednesday, September 28, 2011 15:54:13CBH is now operational again. Bias settings are 1.3V at 20mA, "1st stage" only. The other stage bias controllers are not used.
crioscrios@naic.eduThursday, February 2, 2012 10:01:06The receiver started cooldonw on Jan. 25th 2012. The sys is connected with 430Mhz, CBL.
Denis Urbaindurbain@naic.eduThursday, October 11, 2012 08:24:06On the 3rd of Oct, the cold head was stopped due to contamination. On the 4th the receiver was warm and the cold head was removed, water was found inside the canister. The displacer was left overnight in the oven under vacuum to dry. On the 5th the cold head was installed and purged and the vacuum pump was connected and the in the afternoon we started cooling.The vacuum pump was stopped the next day.
Denis Urbaindurbain@naic.eduFriday, May 24, 2013 16:54:18The C band high was removed in May form the rotary floor to make room for the MIT antenna and compensate for the weight added by the MIT equipment. The cold-head was cleaned and the displacer baked in the oven. The receiver was installed back at the end of the experiment.
Denis Urbaindurbain@naic.eduWednesday, September 25, 2013 09:36:59The receiver had to be removed from the rotary floor during the AO19 experiment. It was then re-installed at the end of the campaign.
Lost Vacuum felix.fernandez@naic.eduTuesday, April 22, 2014 10:15:45CBH's temperature started to increase on 21-APR-2014 at 15:00. After today's inspection we found that the compressor had lost most of its He. The pressure readings were 10/100 (return/supply in PSI). Furthermore, CBH vacuum was about 100 mTorr. We decided to let CBH warm up and then reestablish vacuum and finally cool-down. As for the compressor, we'll service it after today's s-radar run.
Compressor failurefelix.fernandez@naic.eduWednesday, July 30, 2014 11:10:29The SBN, SBL/W , and CBH compressor (model M600) went "dead" after last Jun-19-2014 power failure. After several attempts to fix the problem while the compressor was upstairs it was concluded that the compressor's scroll pump was seized and was tripping the over-current sensor. Hence, SBN, SBL/W and CBH have been warm since that event. Today we finally got the opportunity to replace the compressor. All the aforementioned receivers will begin their vacuum and cool-down cycles sequentially today.
Cold-head contaminationfelix.fernandez@naic.eduFriday, August 1, 2014 10:56:32After the cool-down process ended the cold-head started ratcheting. We will warmup and decontaminate the cold-head.
Started Cool-down Cyclefelix.fernandez@naic.eduMonday, August 4, 2014 13:48:47After checking the cold-head we did not find any remains of moisture. Most likely the contamination came from some foreign gas. Hence, the cold-head was flushed and replaced. We started the cool-down cycle today.
Bad tsys valuescrios@naic.eduMonday, October 6, 2014 14:35:03LNA were found off . LNA turned back on Tsys values normal.
Cold-head contaminationfelix.fernandez@naic.eduMonday, November 10, 2014 10:55:05Last week CBH's cold-head was knocking. The cold-head was taken out and its displacers were baked to remove humidity contamination. We started the cool-down process remotely assuming a good enough vacuum (we did not physically check we just allowed the vacuuming process to run for a whole day). Last Friday CBH's temperature started to climb and today we found out that CBH's temperature stabilized around 200K over the weekend. During today's inspection we found out that CBH had lost its vacuum. Hence, we restarted the thermal cycle once again. This time around we will physically check the vacuum before starting the cool-down process.
CBH warmed upfelix.fernandez@naic.eduFriday, January 8, 2016 09:27:31During the last weeks of Decembre 2015 the CBH Rx started to warm up at a slow rate. During Jan-3-2016 evening the temperature started to climb at a fast pace. On Monday Jan-4 the CBH had reached almost 300K. We concluded that CBH had slowly lost its vacuum. We will attend CBH once the WD campaign is over.


<> Back to Top